OpenStudy (anonymous):

solve for x: -3|2x+6| = -12 this is about absolute values A. x=1 & x= 5 B. x= -1 and x= -5 C. x= -9 and x= 3 D. No solution

OpenStudy (jdoe0001):

\( { -3|2x+6| = -12\implies |2x+6| = \cfrac{-12}{-3}\implies \begin{cases} +(2x+6)=\frac{-12}{-3}&x=?\\ -(2x+6)=\frac{-12}{-3}&x=? \end{cases} }\)

OpenStudy (anonymous):

im confused, @jdoe0001

OpenStudy (jdoe0001):

is an absolute value expression thus "x" will take on two values one for each scenario

OpenStudy (anonymous):

for this i believe there is 2 answers